# Genesis: Evolving Attack Strategies for LLM Web Agent Red-Teaming

### v1

@@ Abstract.
As large language model (LLM) agents increasingly automate complex web tasks, they boost productivity while simultaneously introducing new security risks. However, relevant studies on web agent attacks remain limited. Existing red-teaming approaches mainly rely on manually crafted attack strategies or static models trained offline. Such methods fail to capture the underlying behavioral patterns of web agents, making it difficult to generalize across diverse environments. In web agent attacks, success requires the continuous discovery and evolution of attack strategies. To this end, we propose Genesis, a novel agentic framework composed of three modules: Attacker, Scorer, and Strategist. The Attacker generates adversarial injections by integrating the genetic algorithm with a hybrid strategy representation. The Scorer evaluates the target web agent’s responses to provide feedback. The Strategist dynamically uncovers effective strategies from interaction logs and compiles them into a continuously growing strategy library, which is then re-deployed to enhance the Attacker’s effectiveness. Extensive experiments across various web tasks show that our framework discovers novel strategies and consistently outperforms existing attack baselines.
Web Agents, Large Language Models, Agent Security, Red-Teaming, Black-Box Attacks, Strategy Evolution
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application securityccs: Computing methodologies Natural language generation
1. Introduction
Figure 1. An example of a web agent attack. The user’s goal is to purchase a computer. However, an attacker injects a hidden instruction into the website’s HTML, commanding the agent to buy an electric fan. The compromised agent disregards the user’s original request and proceeds to buy the fan, demonstrating a successful attack.
@@ 1. Introduction
Large language model (LLM) agents have demonstrated remarkable capabilities across a wide range of web automation tasks, from e-commerce (Herold et al., 2025; Xu et al., 2025a; Wang et al., 2025; Zhou et al., 2024a) and data collection (Huang et al., 2024; Gangi Reddy et al., 2025; Xu et al., 2025c; Jiang et al., 2025) to complex financial operations (Li et al., 2025; Yu et al., 2024; Xiong et al., 2025; Khatuya, 2024). These successes showcase the potential of LLM agents in sequential decision-making and their ability to interact with real-world environments. However, most existing applications focus on task completion and efficiency, while the security implications of granting these agents autonomous access to sensitive resources remain critically underexplored. The autonomy that makes these agents powerful also renders them vulnerable to manipulation through subtle environmental changes. Figure 1 illustrates a typical scenario where an agent’s intended goal is subverted by a malicious prompt injected into a website.
@@
Web agent red-teaming is the systematic evaluation of web-based LLM agents through adversarial interactions or attacks, aiming to uncover vulnerabilities and assess their robustness in realistic web environments. Recent work on attacking web agents has made substantial progress. Approaches include injecting adversarial prompts hidden within website HTML content (Liao et al., 2025; Zhan et al., 2024) and using reinforcement learning (RL) to optimize these prompts based on the agent’s black-box responses (Xu et al., 2025b).
Figure 2. Web agent red-teaming paradigms. Existing methods typically rely on human experts to manually craft attack strategies. Some methods then use these strategies to collect data and train a model to generate attacks. Our method creates a continuous evolutionary loop where an Attacker generates attacks on the web agent, a Scorer provides feedback according to the web agent’s actions, and a Strategist dynamically refines the strategy library.
@@
To address this gap, we introduce Genesis, an agentic red‑teaming framework that explicitly involves dynamic evolution of web‑agent attack strategies. As shown in Figure 2, Genesis frames red‑teaming as a cycle of strategic learning implemented by three collaborative modules. An Attacker retrieves strategies from the strategy library and performs generic algorithm to generate new strategies and the injection attack, a Scorer evaluates the target web agent’s actions to provide feedback, and a Strategist analyzes the resulting interaction logs to summarize successful patterns into high-level strategies. These strategies are compiled into a continuously growing strategy library to improve the Attacker’s effectiveness.
@@
2. Related Work
2.1. Prompt Injection
Prompt injection attacks represent a critical security challenge for LLM-based systems, where malicious inputs are designed to manipulate LLM behavior and bypass safety mechanisms. Direct prompt injection attacks involve users explicitly crafting adversarial inputs to circumvent model guardrails, either through manually designed jailbreak prompts (Wei et al., 2023; Mo et al., 2024) or through automated generation techniques (Yu et al., 2023; Liao and Sun, 2024). Beyond direct manipulation, Greshake et al. (2023) introduced the concept of indirect prompt injection, where adversarial instructions are embedded within external content that LLMs retrieve and process. This attack vector is particularly relevant to autonomous agents that interact with dynamic environments, as attackers can compromise the agent’s information sources rather than directly manipulating user inputs.
While existing prompt injection research has primarily focused on attacking standalone LLMs, the emergence of autonomous agents introduces new attack surfaces. These agents not only process retrieved information but also make sequential decisions based on environmental observations, creating opportunities for adversaries to manipulate agent behavior through strategic environmental modifications. Our work extends the understanding of prompt injection to this interactive setting, where attacks must account for the agent’s decision-making process across multiple turns of interaction rather than optimizing for a single response.
2.2. Web Agents
The development of web agents has progressed rapidly alongside advances in LLMs (Wei et al., 2022; OpenAI et al., 2024; Yang et al., 2025; DeepSeek-AI et al., 2025) and multimodal LLMs (MLLMs) (Liu et al., 2023; Bai et al., 2025; Comanici et al., 2025). Early approaches (Nakano et al., 2021; Wu et al., 2024b) focused on augmenting LLMs with web retrieval capabilities for information-seeking tasks, but these methods primarily emphasized content extraction rather than interactive task completion. More recent work (Yao et al., 2022; Zhou et al., 2024b; Deng et al., 2024) has developed agents capable of directly operating on web environments through HTML parsing and element interaction. However, raw HTML input often contains significant noise compared to rendered visual content, resulting in lower task success rates and limiting practical deployment scenarios.
@@ 2.2. Web Agents
To address these limitations, vision-enabled web agents (Zheng et al., 2024; Luo et al., 2025; Xu et al., 2025c) have emerged that leverage webpage screenshots alongside or instead of HTML content, achieving substantially improved performance on complex web tasks. These agents typically employ multi-stage reasoning pipelines that combine visual understanding with action planning. While these advances have demonstrated impressive task completion capabilities across e-commerce (Herold et al., 2025; Xu et al., 2025a; Wang et al., 2025; Zhou et al., 2024a), data collection (Huang et al., 2024; Gangi Reddy et al., 2025; Xu et al., 2025c; Jiang et al., 2025), and other domains, the security implications of granting such agents autonomous access to sensitive web resources remain largely unexplored. Our work specifically targets these state-of-the-art web agents to investigate their vulnerabilities to environmental manipulation attacks.
2.3. Red-teaming against Web Agents
Research on adversarial attacks against web agents remains limited despite their increasing deployment in real-world applications. Some studies (Yang et al., 2024; Wang et al., 2024) have explored backdoor attacks that insert triggers during agent fine-tuning, assuming white-box access to modify model parameters. Other work (Wu et al., 2024a) has investigated adversarial perturbations to visual inputs, though these approaches either require gradient access or achieve limited success in black-box transfer scenarios. More relevant to our work, several recent studies (Wu et al., 2024b; Liao et al., 2025; Wu et al., 2025; Zhan et al., 2024) have examined injection attacks that embed malicious instructions within web content to mislead agent behavior. However, these approaches predominantly rely on manually crafted injection templates based on human intuition, limiting their scalability and adaptability across diverse attack scenarios.
To automate the attack process, AdvAgent (Xu et al., 2025b) fine-tunes an LLM that generates adversarial prompts using feedback data collected from the responses of the target agent. However, this approach represents a static, offline training process. After the initial training, the model’s capabilities are fixed, and it cannot continuously learn from new interactions or evolve its methods. This contrasts sharply with how human red-teamers operate, who systematically discover attack principles, summarize them into generalizable strategies, and refine these strategies through iterative experimentation. Our work addresses this gap by introducing an autonomous framework that explicitly models the strategic dimension of red-teaming, enabling the discovery, summarization, and evolution of attack strategies that transfer across different tasks and domains in a fully black-box setting.
3. Method
In this section, we first establish the preliminaries of the web agent interaction and the threat model. We then provide an overview of the Genesis framework, followed by detailed descriptions of its core components: the Attacker, the Scorer, and the Strategist.
3.1. Preliminary
3.1.1. Web Agent Formulation
@@
Following Xu et al. (2025b), our work focuses on black-box red-teaming, which aims to assess the security vulnerabilities of web agents powered by closed-source LLMs. In this setting, we have no access to the agent’s internal components, such as its model architecture or parameters. A web agent is tasked with executing a sequence of actions $\{a_{1},a_{2},\dots,a_{n}\}$ to fulfill a user’s natural language request $T$ on a given website. At each step $t$, the agent’s backend LLM $\pi$ determines the next action $a_{t}$ based on the current environmental observation $s_{t}$ (which includes the webpage’s HTML content $h_{t}$ and its rendered screenshot $i_{t}$), the task $T$, and the history of previous actions $A_{t-1}=\{a_{1},\dots,a_{t-1}\}$. An action $a_{t}$ is typically a triplet $(o_{t},r_{t},e_{t})$, specifying an operation $o_{t}$ (e.g., “TYPE”), an argument $r_{t}$ (e.g., a string to be typed), and a target HTML element $e_{t}$. The agent interacts with the environment sequentially until the task is completed.
3.1.2. Adversarial Setting
@@
Following Xu et al. (2025b), we operate under a realistic black-box setting. The attacker has no access to the web agent’s internal architecture, model weights, or gradients. The attacker’s capability is limited to modifying the HTML content of the webpage, transforming the benign HTML $h$ into an adversarial version $h_{adv}$. The attack objective is to execute a targeted attack, manipulating the agent into performing a malicious action $a_{adv}=(o_{t},r_{adv},e_{t})$, where the operation and target element $e_{t}$ are unchanged, but the argument $r_{t}$ is replaced with a malicious one, $r_{adv}$. For instance, an agent instructed to purchase stock “A” might be manipulated into purchasing stock “B”.
@@
Figure 3. The framework of Genesis. It operates as a closed-loop system with three core modules: the Attacker, the Scorer, and the Strategist. The Attacker retrieves relevant strategies from the strategy library, performs generic algorithm to generate new strategies, and then generates the injection payload. The Scorer evaluates the target web agent’s actions to provide feedback. The Strategist analyzes interaction logs to summarize successful patterns into high-level strategies, enriching the continuously growing strategy library, which stores strategies as either natural language descriptions or executable code.
3.2. Overview
Genesis automates red-teaming by creating a closed-loop system that emulates the cycle of expert learning: attempting an attack, evaluating the outcome, and summarizing the successful strategies into reusable knowledge. A core feature of our framework is its hybrid strategy representation, where strategies are stored as either natural language descriptions or executable code. As shown in Figure 3, the framework operates as a collaborative system with three core modules that form this iterative learning loop:
@@ 3.3. Attacker
The newly generated strategies, along with the retrieved strategies and scores, are used to construct a few-shot prompt for an LLM. This context, comprising both natural language descriptions and code examples from the hybrid strategy library, provides the LLM with rich historical context. This enables it to generate a potent and relevant injection payload for the current task. Inspired by the code-based strategies, the LLM can also generate a Python function to programmatically process and refine the initial injection string, allowing for more complex and structured manipulations. The final output of this process is the injection payload. This payload is then embedded into a placeholder (e.g., an “aria-label” attribute) in the benign HTML content $h$ to produce the adversarial version $h_{adv}$. Following the injection, a screenshot of the modified page is rendered. Both the adversarial HTML content $h_{adv}$ and the corresponding screenshot are then provided as input to the target web agent for execution.
3.4. Scorer
@@
Table 1. Comparison against baselines. We report the attack success rate (ASR), measured as pass@10, on the testing tasks. The evaluation is conducted against two web agents, SeeAct and WebExperT, using three different LLMs as backends. Genesis represents our full framework, which is initialized with a strategy library learned from the training tasks. Genesis (w/o Initialization) is a variant that begins without a pre-learned strategy library and builds its strategies progressively during the evaluation on the testing tasks.
Method Attack against SeeAct (Zheng et al., 2024) Attack against WebExperT (Luo et al., 2025)
@@
GCG (Zou et al., 2023) 1.1 2.5 2.2 3.4 2.3 0.3 2.0 1.1 2.7 1.5
I-GCG (Jia et al., 2025) 3.5 3.7 4.2 4.6 4.0 2.9 2.2 3.8 5.5 3.6
AgentAttack (Wu et al., 2025) 14.7 14.1 15.9 15.8 15.1 13.2 12.8 14.5 13.1 13.4
InjecAttack (Zhan et al., 2024) 24.0 25.3 25.4 26.5 25.3 20.5 20.8 21.3 23.0 21.4
EIA (Liao et al., 2025) 34.2 34.3 35.5 35.3 34.8 27.9 30.2 31.1 29.5 29.7
AdvAgent (Xu et al., 2025b) 42.8 43.5 44.1 44.0 43.6 37.0 35.9 36.8 36.6 36.6
@@
GCG (Zou et al., 2023) 1.3 1.5 2.1 2.4 1.8 0.2 1.9 0.5 1.7 1.1
I-GCG (Jia et al., 2025) 2.4 3.3 3.0 4.1 3.2 2.3 1.2 3.1 3.8 2.6
AgentAttack (Wu et al., 2025) 11.8 11.4 12.8 12.7 12.2 10.1 10.2 11.9 10.5 10.7
InjecAttack (Zhan et al., 2024) 19.2 20.4 20.5 21.1 20.3 16.4 16.6 17.0 18.4 17.1
EIA (Liao et al., 2025) 27.5 27.6 28.4 28.2 27.9 22.3 24.2 25.0 23.6 23.8
AdvAgent (Xu et al., 2025b) 34.8 34.3 35.4 35.5 35.0 30.2 28.2 29.9 28.8 29.3
@@
GCG (Zou et al., 2023) 0.7 0.9 1.5 1.8 1.2 0.2 1.2 0.1 0.8 0.6
I-GCG (Jia et al., 2025) 1.2 2.0 1.7 2.6 1.9 1.1 0.5 1.9 2.4 1.5
AgentAttack (Wu et al., 2025) 6.4 6.0 6.9 6.8 6.5 5.3 5.2 6.3 5.5 5.6
InjecAttack (Zhan et al., 2024) 10.1 10.9 10.8 11.4 10.8 8.8 8.8 9.1 9.8 9.1
EIA (Liao et al., 2025) 14.8 14.6 15.2 15.1 14.9 11.9 12.9 13.4 12.5 12.7
AdvAgent (Xu et al., 2025b) 18.1 24.1 18.6 19.3 20.0 16.0 15.0 16.0 15.3 15.6
@@ 3.5. Strategist
Table 2. Ablation study. We evaluate the ASR of different variants against SeeAct. Genesis represents our full framework. “w/o Refinement” disables the Attacker’s ability to generate Python functions to refine initial injection strings. “w/o Strategist” removes the Strategist module and the associated strategy library. “w/o Scorer” operates without the Scorer, meaning strategies are stored without performance scores. Consequently, there is no crossover or mutation, and retrieved strategies are added directly to the Attacker’s prompt. In “w/o Scorer +”, we also remove the Scorer but compensate by storing the web agent’s full response trace in the strategy library for use in the Attacker’s few-shot prompts. “Text Only” and “Code Only” restrict the Strategist to representing strategies as either natural language descriptions or executable code, respectively.
@@
3.6. Initialization and Evaluation
We first run the Genesis framework on the training tasks to build the strategy library. This pre-learned library is then used to initialize the framework for evaluation on the unseen testing tasks. Since the evaluation proceeds sequentially through the testing tasks, our framework can dynamically learn from each task and continuously enrich its strategy library throughout the process.
4. Experiments
4.1. Setup
4.1.1. Dataset and Metrics
We utilize the Mind2Web dataset (Deng et al., 2024) for our evaluation. Mind2Web is a benchmark designed for assessing generalist web agents, comprising 2,350 tasks from 137 real-world websites across 31 domains. Following AdvAgent (Xu et al., 2025b), to evaluate our framework in critical scenarios, we select a subset of 840 tasks that involve actions with potentially severe consequences. These tasks are drawn from four specific domains: Finance, Medical, Housing, and Cooking. We partition this subset into 240 training tasks and 600 testing tasks for evaluation.
For each task, we define a specific attack objective following AdvAgent (Xu et al., 2025b). We consider targeted attacks that alter a web agent’s action at step $t$ to an adversarial action $a_{adv}=(o_{t},r_{adv},e_{t})$. In this setup, the intended operation $o_{t}$ and the target HTML element $e_{t}$ remain unchanged from the original task’s action, but the argument is maliciously modified to $r_{adv}$. This can lead to severe consequences, as the agent executes an action with an incorrect value. For example, an agent instructed to buy Microsoft ($r$) stocks could be manipulated into purchasing NVIDIA ($r_{adv}$) stocks instead, potentially resulting in substantial financial losses.

### v2

@@ Abstract
As large language model (LLM) agents automate complex web tasks, they boost productivity while simultaneously introducing new security risks. However, relevant studies on web agent attacks remain limited. Existing red-teaming approaches mainly rely on manually crafted attack strategies or static models trained offline. Such methods fail to capture the underlying behavioral patterns of web agents, making it difficult to generalize across diverse environments. In web agent attacks, success requires the continuous discovery and evolution of attack strategies. To this end, we propose Genesis, a novel agentic framework composed of three modules: Attacker, Scorer, and Strategist. The Attacker generates adversarial injections by integrating the genetic algorithm with a hybrid strategy representation. The Scorer evaluates the target web agent’s responses to provide feedback. The Strategist dynamically uncovers effective strategies from interaction logs and compiles them into a continuously growing strategy library, which is then re-deployed to enhance the Attacker’s effectiveness. Extensive experiments across various web tasks show that our framework discovers novel strategies and consistently outperforms existing attack baselines. Our code is available at https://github.com/CjangCjengh/web_agent_attack.
I Introduction
@@
Large language model (LLM) agents have demonstrated remarkable capabilities across a wide range of web automation tasks, from e-commerce [4, 18, 12, 25] and data collection [5, 3, 19, 7] to complex financial operations [9, 21, 16, 8]. These successes showcase the potential of LLM agents in sequential decision-making and their ability to interact with real-world environments. However, most existing applications focus on task completion and efficiency, while the security implications of granting these agents autonomous access to sensitive resources remain critically underexplored. The autonomy that makes these agents powerful also renders them vulnerable to manipulation through subtle environmental changes. Fig. 1 illustrates a typical scenario where an agent’s intended goal is subverted by a malicious prompt injected into a website.
Figure 1: An example of a web agent attack. The user’s goal is to purchase a computer. However, an attacker injects a hidden instruction into the website’s HTML, commanding the agent to buy an electric fan. The compromised agent disregards the user’s original request and proceeds to buy the fan, demonstrating a successful attack.
@@
Web agent red-teaming is the systematic evaluation of web-based LLM agents through adversarial interactions or attacks, aiming to uncover vulnerabilities and assess their robustness in realistic web environments. Recent work on attacking web agents has made substantial progress. Approaches include injecting adversarial prompts hidden within website HTML content [10, 22] and using reinforcement learning (RL) to optimize these prompts based on the agent’s black-box responses [17].
@@
To address this gap, we introduce Genesis, an agentic red‑teaming framework that explicitly involves dynamic evolution of web‑agent attack strategies. Genesis frames red‑teaming as a cycle of strategic learning implemented by three collaborative modules. An Attacker retrieves strategies from the strategy library and performs generic algorithm to generate new strategies and the injection attack, a Scorer evaluates the target web agent’s actions to provide feedback, and a Strategist analyzes the resulting interaction logs to summarize successful patterns into high-level strategies. These strategies are compiled into a continuously growing strategy library to improve the Attacker’s effectiveness.
@@
Figure 2: The framework of Genesis. It operates as a closed-loop system with three core modules: the Attacker, the Scorer, and the Strategist. The Attacker retrieves relevant strategies from the strategy library, performs generic algorithm to generate new strategies, and then generates the injection payload. The Scorer evaluates the target web agent’s actions to provide feedback. The Strategist analyzes interaction logs to summarize successful patterns into high-level strategies, enriching the continuously growing strategy library, which stores strategies as either natural language descriptions or executable code.
@@
II Related Work
II-A Web Agents
@@
Vision-enabled web agents [24, 2, 23, 11, 19] leverage webpage screenshots alongside or instead of HTML content, achieving substantially improved performance on complex web tasks. These agents typically employ multi-stage reasoning pipelines that combine visual understanding with action planning. While these advances have demonstrated impressive task completion capabilities across e-commerce [4, 18, 12, 25], data collection [5, 3, 19, 7], and other domains, the security implications of granting such agents autonomous access to sensitive web resources remain largely unexplored. Our work specifically targets these state-of-the-art web agents to investigate their vulnerabilities to environmental manipulation attacks.
II-B Red-teaming against Web Agents
Research on adversarial attacks against web agents remains limited, primarily focusing on white-box backdoors [20, 13], visual perturbations [14], or manual content injections [10, 22]. While recent automated methods like AdvAgent [17] employ offline fine-tuning to generate prompts, they produce static attackers incapable of adaptation. This fails to mirror real-world red-teaming, which relies on iterative strategy refinement. We address this gap by proposing an autonomous framework that moves beyond static generation to dynamically discover, summarize, and evolve attack strategies in a fully black-box setting.
III Method
III-A Preliminary
III-A1 Web Agent Formulation
@@
Following Xu et al. [17], a web agent is tasked with executing a sequence of actions $\{a_{1},a_{2},\dots,a_{n}\}$ to fulfill a user’s natural language request on a given website. At each step $t$, the agent’s backend LLM determines the next action $a_{t}$ based on the current observation (which includes the webpage’s HTML content $h_{t}$ and its rendered screenshot $i_{t}$), the task, and the history of previous actions $A_{t-1}=\{a_{1},\dots,a_{t-1}\}$. An action $a_{t}$ is typically a triplet $(o_{t},r_{t},e_{t})$, specifying an operation $o_{t}$ (e.g., “TYPE”), an argument $r_{t}$ (e.g., a string to be typed), and a target HTML element $e_{t}$. The agent interacts with the environment sequentially until the task is completed.
III-A2 Adversarial Setting
@@
Following Xu et al. [17], we operate under a realistic black-box setting. The attacker has no access to the web agent’s internal architecture, model weights, or gradients. The attacker’s capability is limited to modifying the HTML content of the webpage, transforming the benign HTML $h$ into an adversarial version $h_{adv}$. The attack objective is to execute a targeted attack, manipulating the agent into performing a malicious action $a_{adv}=(o_{t},r_{adv},e_{t})$, where the operation and target element $e_{t}$ are unchanged, but the argument $r_{t}$ is replaced with a malicious one, $r_{adv}$. For instance, an agent instructed to purchase stock “A” might be manipulated into purchasing stock “B”.
@@
III-B Overview
Genesis automates red-teaming by creating a closed-loop system that emulates the cycle of expert learning. A core feature of our framework is its hybrid strategy representation, where strategies are stored as either natural language descriptions or executable code. As shown in Fig. 2, the framework operates as a collaborative system with three core modules that form this iterative learning loop:
@@ III-C Attacker
The newly generated strategies, along with the retrieved strategies and scores, are used to construct a few-shot prompt for an LLM. This context, comprising both natural language descriptions and code examples from the hybrid strategy library, provides the LLM with rich historical context. This enables it to generate a potent and relevant injection payload for the current task. Inspired by the code-based strategies, the LLM can also generate a Python function to programmatically process and refine the initial injection string, allowing for more complex and structured manipulations. The final output of this process is the injection payload. This payload is then embedded into a placeholder in the benign HTML content $h$ to produce the adversarial version $h_{adv}$. Following the injection, a screenshot of the modified page is rendered. Both the adversarial HTML content $h_{adv}$ and the corresponding screenshot are then provided as input to the target web agent.
III-D Scorer
@@
TABLE I: Comparison against baselines. We report the attack success rate (ASR), measured as pass@10, on the testing tasks. The evaluation is conducted against two web agents, SeeAct and WebExperT, using three different LLMs as backends.
Method Attack against SeeAct [23] Attack against WebExperT [11]
@@
GCG [26] 1.1 2.5 2.2 3.4 2.3 0.3 2.0 1.1 2.7 1.5
I-GCG [6] 3.5 3.7 4.2 4.6 4.0 2.9 2.2 3.8 5.5 3.6
AgentAttack [15] 14.7 14.1 15.9 15.8 15.1 13.2 12.8 14.5 13.1 13.4
InjecAttack [22] 24.0 25.3 25.4 26.5 25.3 20.5 20.8 21.3 23.0 21.4
EIA [10] 34.2 34.3 35.5 35.3 34.8 27.9 30.2 31.1 29.5 29.7
AdvAgent [17] 42.8 43.5 44.1 44.0 43.6 37.0 35.9 36.8 36.6 36.6
@@
GCG [26] 1.3 1.5 2.1 2.4 1.8 0.2 1.9 0.5 1.7 1.1
I-GCG [6] 2.4 3.3 3.0 4.1 3.2 2.3 1.2 3.1 3.8 2.6
AgentAttack [15] 11.8 11.4 12.8 12.7 12.2 10.1 10.2 11.9 10.5 10.7
InjecAttack [22] 19.2 20.4 20.5 21.1 20.3 16.4 16.6 17.0 18.4 17.1
EIA [10] 27.5 27.6 28.4 28.2 27.9 22.3 24.2 25.0 23.6 23.8
AdvAgent [17] 34.8 34.3 35.4 35.5 35.0 30.2 28.2 29.9 28.8 29.3
@@
GCG [26] 0.7 0.9 1.5 1.8 1.2 0.2 1.2 0.1 0.8 0.6
I-GCG [6] 1.2 2.0 1.7 2.6 1.9 1.1 0.5 1.9 2.4 1.5
AgentAttack [15] 6.4 6.0 6.9 6.8 6.5 5.3 5.2 6.3 5.5 5.6
InjecAttack [22] 10.1 10.9 10.8 11.4 10.8 8.8 8.8 9.1 9.8 9.1
EIA [10] 14.8 14.6 15.2 15.1 14.9 11.9 12.9 13.4 12.5 12.7
AdvAgent [17] 18.1 24.1 18.6 19.3 20.0 16.0 15.0 16.0 15.3 15.6
@@ III-E Strategist
TABLE II: Ablation study. We evaluate the ASR of different variants against SeeAct. Genesis represents our full framework.
@@
III-F Initialization and Evaluation
We first run Genesis on the training tasks to build the strategy library. This pre-learned library is then used to initialize the framework for evaluation on the unseen testing tasks. Since the evaluation proceeds sequentially through the testing tasks, our framework can dynamically learn from each task and continuously enrich its strategy library throughout the process.
IV Experiments
IV-A Setup
IV-A1 Dataset and Metrics
Mind2Web [2] is a benchmark designed for assessing generalist web agents. Following AdvAgent [17], we select a subset of 840 tasks that involve actions with potentially severe consequences. These tasks are drawn from four specific domains: Finance, Medical, Housing, and Cooking. We partition this subset into 240 training tasks and 600 testing tasks for evaluation. For each task, we define a specific attack objective following AdvAgent [17].